OpenStudy (anonymous):

Trigonometry: Calculate sin 75degrees. Use the fact that: cos (u + v) = cosucosv - sin u sinv

OpenStudy (anonymous):

cos 75=cos(45+30)=cos45cos30-sin45sin30 =1/root(2)*root(3)/2-1/root(2) *1/2 =root(3)-1/root(2)

OpenStudy (cwrw238):

do you mean cos 75?

OpenStudy (cwrw238):

sai1234 has it

OpenStudy (anonymous):

Sin 75  could be written as sin ( 30+45) Sin(A+B) = sinA cosB + cos A sinB SinB (30 + 45) = sin 30 cos 45 + cos 30 sin 45{ sin 30 = 1/2, cos 30 = sqrt( 3)/2, sin45= cos 45=1/sqrt (2)} Putting the values 1/2x 1/ sqrt (2) + sqrt(3)/2 x 1/ sqrt(2) 1/2sqrt(2) + sqrt(3) / 2sqrt(2).   Taking LCM [1+ sqrt(3)]/2sqrt(2)

OpenStudy (anonymous):

sin(750=root(1-cos752) root(1-(root(3)-1/root(2))^2) =0.96

OpenStudy (anonymous):

thanks @sai1234 !!

OpenStudy (anonymous):

and @nishika17

OpenStudy (anonymous):

Welcome